OpenStudy (anonymous):

If someone could make sure that my work is correct on this, I would be greatful: Find the accumulated amount A if the principle P=$11,000 is invested at the rate of r=5% per year for t=5.5 years, compounded quarterly. My work can be seen below...

OpenStudy (anonymous):

ok, i used i=r/m formula to get i=.05/4=.0125 per qtr. 4 qtrs per year x 5.5 years gives me 22 total qtrs. Plug that into A=P(1+rt)^n to get A=11000(1+.0125)^22 A= $14,457.17